OpenStudy (anonymous):

Express in terms of sums and differences of logarithms. logcubedrootxy^2 please help

OpenStudy (anonymous):

solve\[\log2(x-3)+\log2(2x-10)=4\]

jimthompson5910 (jim_thompson5910):

Is the problem \[\Large \log\left(\sqrt[3]{x}y^2\right)\] or Is the problem \[\Large \log\left(\sqrt[3]{xy^2}\right)\]

OpenStudy (anonymous):

the second one. Yeah sorry i couldnt figure out how to type it this is the first time iv used this website. lol

jimthompson5910 (jim_thompson5910):

\[\Large \log\left(\sqrt[3]{xy^2}\right)\] \[\Large \log\left((xy^2)^{\frac{1}{3}}\right)\] \[\Large \frac{1}{3}\log\left(xy^2\right)\] \[\Large \frac{1}{3}\left(\log(x) + \log(y^2)\right)\] \[\Large \frac{1}{3}\left(\log(x) + 2\log(y)\right)\] \[\Large \frac{1}{3}\log(x) + \frac{1}{3}*2\log(y)\] \[\Large \frac{1}{3}\log(x) + \frac{2}{3}\log(y)\]

OpenStudy (anonymous):

THANK YOU SO MUCH

jimthompson5910 (jim_thompson5910):

yw